OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131 Please explain how you got your answer.

OpenStudy (anonymous):

what are you adding each time?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

right so maybe you can use something like \[\sum_{k=0}^n -4+9k\]

OpenStudy (anonymous):

course you still got to figure out what \(n\) is

OpenStudy (anonymous):

Im assuming it's infinity considering the other choice I have is 15 and the 15th term would not be 131

OpenStudy (anonymous):

\[-4+9n=131\] what is \(n\)?

OpenStudy (anonymous):

oh... 15....

OpenStudy (anonymous):

ahah. Thanks!

OpenStudy (anonymous):

ok so if \[\sum_{k=0}^{15} -4+9k\] is a choice, pick that one

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

It is. Mind helping me with another?

OpenStudy (anonymous):

go ahead and ask, maybe i can helpl

OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 16 + 25 + 36 + 49 + ... + n^2 + ...

OpenStudy (anonymous):

does it really have \(...\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then just go with \[\sum_{n=4}^{\infty}n^2\]

OpenStudy (anonymous):

silly because you cannot add all this stuff up but whatever

OpenStudy (anonymous):

Why n=4 at the bottom?

OpenStudy (anonymous):

what number starts the summation ?

OpenStudy (anonymous):

Don't know. Would've said 16.

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

and what number squared gives 16?

OpenStudy (anonymous):

4. So It's the square root of the first term that goes on the bottom???

OpenStudy (anonymous):

\[\sum_{n=4}^{\infty}n^2=4^2+5^2+6^2+...\]

OpenStudy (anonymous):

each term is the square of some number

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

starting with 16, that is the square of 4

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

yw again